OpenStudy (anonymous):

a larger cone with a slant of 42 yards. and a smaller cone with a slant of 30 yards and radius of 9 yards. What is the approximate surface area of the larger cone? Answer 978.5 square yards 1,103.1 square yards 1544.4 square yards 2,162.3 square yards

OpenStudy (anonymous):

?

OpenStudy (callisto):

slant of larger cone radius of larger cone ----------------- = ------------------ slant of smaller cone radius of smaller cone 42 radius of larger cone --- = ------------------ 30 9 radius of larger cone = 42/30 x 9 = 12.6 surface area = πrl + πr^2 = π (12.6)(42) + π(12.6)^2 Can you do calculation ? Do you understand the workings?

OpenStudy (anonymous):

would it be 2,162.3 ?

OpenStudy (callisto):

I got 2161.3 :S

OpenStudy (anonymous):

close enough lol thanks

OpenStudy (callisto):

You're welcome, I'm confused with my answer :(